OpenStudy (anonymous):

if 2 sides of a rectangle are 7x-3 yards, and the other 2 are 7x+3 yards then what is the area?

OpenStudy (bibby):

rectangle area = l*w

OpenStudy (anonymous):

i know dat :/

OpenStudy (anonymous):

but wat is the lenght times the width in this particular equation?

OpenStudy (bibby):



OpenStudy (anonymous):

i know that, but i don't know how to figure it out with all of the x's

OpenStudy (bibby):

multiply term by term \((7x-3)(7x+3)=(7x*7x)+(7x*3)-(7x*3)-(3*3)\)

OpenStudy (bibby):

or just remember that \((a-b)(a+b)=a^b-b^2\)

OpenStudy (bibby):

\((a-b)(a+b)=a^2-b^2\) *

OpenStudy (anonymous):

so that would mean that the answer is what?

OpenStudy (funnyguy2):

hi

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

(7x+3)*(7x-3)=49x^2-9 (a-b)*(a+b)=a^2-b^2

OpenStudy (anonymous):

ok thanks :)